Tex co issued bonds at the price of 0.97, which is at face value $ 200,000,000. This version will affect the accounting equation to:

Group of answer choices

Assets and liabilities will both increase by $ 194 million.

Assets will increase by $ 200,000,000 and liabilities will increase by $ 194 million.

Assets and liabilities will both increase by $ 200,000,000.

Assets will increase by $ 194,000,000 and debt will increase by $ 200,000,000.
Answer #1

Answer

--Correct Answer = Option #1: Assets and Liabilities will both increase by $ 194 millions.

>Asset will increase by the amount of Cash received = $ 200 million x 97/100 = $ 194 million
>Liabilities will increase by face value - discount on issuance of bonds = 200 million - (200 - 194) = 200 - 6 = $ 194 millions.